OpenStudy (anonymous):

Need help with this 2 calculus problems I'll post the pic in a sec. C: Please and thank you

OpenStudy (anonymous):

Problem 7 &8

OpenStudy (dumbcow):

for avg velocity its basically slope formula using the 2 endpoints -- velocity = change in distance / change in time

OpenStudy (anonymous):

Can you do I) ? So I can see how its done?

OpenStudy (dumbcow):

ok interval [1,3] \[slope = \frac{10.7 - 1.4}{3-1} = \frac{9.3}{2} = 4.65\]

OpenStudy (anonymous):

Ohh I see. Can you do an example for problem 8? And thanks I get problem 7 now

OpenStudy (dumbcow):

ok avg velocity for interval [1,2] \[slope = \frac{f(2) -f(1)}{2-1} = (2\sin 2\pi +3\cos 2\pi)- (2 \sin \pi +3 \cos \pi) = 3-(-3) = 6\]

OpenStudy (anonymous):

Ohh that doesnt seem too hard. And for 7b. Is the answer 0.?

OpenStudy (dumbcow):

hmm not sure...its not zero because you can see that velocity is increasing over time it says to use a graph, is one provided or you can plug the points into a calculator and it will estimate a graph and function for you

OpenStudy (anonymous):

No

OpenStudy (dumbcow):

how have you been estimating instantaneous velocity in class? im drawing a blank, sorry too used to calculus

OpenStudy (anonymous):

They really didnt teach me. They only showed us some graphs that aren't even related to this

OpenStudy (dumbcow):

hmm ok well for #8 you can estimate it by doing a very small interval close to the point --> [.999, 1.001] for t=1

OpenStudy (dumbcow):

but for #7, you dont have an equation to plug in the values only a limited table you have to first estimate a relation or equation to match the table, then use that to evaluate slope for say [2.99, 3.01]

OpenStudy (anonymous):

I got 7aa but not 7b

OpenStudy (dumbcow):

here is screenshot of using excel to make graph and make a trendline to fit data use that equation to estimate instantaneous velocity at t=3

OpenStudy (anonymous):

Just plug in the 3 in the equation

OpenStudy (dumbcow):

no you want velocity...change in distance/ change in time find slope for interval [2.99, 3.01] the smaller the interval the more accurate the instantaneous velocity

OpenStudy (anonymous):

3.56

OpenStudy (dumbcow):

i get something closer to 6

OpenStudy (anonymous):

How?

OpenStudy (dumbcow):

did little rounding \[f(t) = 0.85t^{2} +.9t\] \[f(3.01) = 10.41\] \[f(2.99) = 10.29\] \[slope = \frac{10.41 - 10.29}{3.01-2.99} = 6\]